John owns a bread bakery. When he increases the price of a loaf of bread from R10 to 

R12, the percentage in the quantity demanded is 5%. Calculate the price elasticity of 

demand for bread.


1
Expert's answer
2021-04-23T11:59:23-0400

Initial Price=R10

Increased price=R12

Percentage change in quantity demanded=5%

 

Price elasticity of demand = %change in quantity demanded/%change in prices

 

%change in price= ((Final price-Initial price)/Initial price)*100

"=\frac{(12-10)}{10}\times100"


"=\frac{2}{10}\times100=20%"

 

 

Price elasticity of demand=%change in quantity demanded/%change in prices

 

Price elasticity of demand"=\frac{5}{20}=0.25"

So, the price elasticity of demand for bread is 0.25